# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ V-B1 Vanilla GAN
Advancing the usage of vanilla GAN models, Xiao et al.[162] proposed a federated learning-based framework for distributed GAN-based network traffic generation with self-supervised learning. The approach involves local model training at the edge, where each edge node contains multiple generators, a discriminator, and a classifier. Each generator is trained on a local dataset, while the discriminator helps improve the performance of all generators at the edge. Locally trained discriminators and generators from all edge nodes are then aggregated globally using the FedAVG algorithm. The classifier at each edge provides pseudo-labels, enabling self-supervised training of the GAN models and facilitating autonomous training of the entire system.
@@
In other network applications, vanilla GANs are utilized to generate network traffic data to enhance network intrusion detection, particularly when class imbalance is prevalent in anomalous data. In this effort, Shahriar et al. [133] apply an encoding approach to convert categorical input features into a low-dimensional continuous space. They also employ feature normalization techniques and reduce feature dimensionality using Principal Component Analysis (PCA) to support the GAN model training process.
V-B2 Wasserstein GAN (WGAN)
@@
In  [48], the authors leverage WGAN to obfuscate network traffic, reducing the detection capabilities of various applications. Initially, they derive statistical features for the flow level data they collected followed by normalization across the flows. Then, they use WGAN to collect network traffic data and generate flow-level statistics, followed by applying WGAN again to synthesize network packet traces based on these synthetic flow-level statistics to produce obfuscated traffic. The model includes multiple WGAN modules trained on different types of network traffic based on applications (e.g., Google Chrome, Google Drive), enabling application-level traffic obfuscation.
@@
Researchers have applied the idea of converting network traffic data into the image domain, where WGANs show great promise, after making multiple modifications to improve GANs for 1D data generation. In [137], the authors employ the Gramian Angular Summation Field (GASF) imaging mechanism to transform certain attributes of network traces into visual representations [158]. These images are utilized for network traffic generation through a simple WGAN framework. This imaging approach effectively captures the correlations between time series points, which previous studies attempt to model using more complex GAN architectures incorporating LSTMs and other networks. The authors develop a post-processing mechanism for the generated images and evaluate the model across multiple traffic types, demonstrating improved performance with the straightforward WGAN for image generation.
The use of WGANs also extends to the generation of radio frequency (RF) signals. However, training GANs on long time-domain sequences of RF signals presents significant difficulties. To address this,  [130] proposed representing RF signals in the spectral domain, allowing the use of existing GAN architectures. The authors consider 2.5s long 4G LTE signal samples at 40 MSamples/sec which were transformed into 256x256 compact spectrograms, where the horizontal axis represented the time dimension over approximately 2 ms, and the vertical axis represented the frequency offset within a 40 MHz band. Each spectrogram served as input to the GAN, effectively framing the RF signal generation as a 2D problem similar to image synthesis. The authors also experimented with direct training of GANs on time series data. They found that while this approach could generate raw I/Q samples, it was significantly more challenging to capture the fine-grained features and multiscale details of LTE signals compared to the spectral domain representation.
V-B3 Conditional GAN (CGAN)
@@
V-B4 GANs with sequential models
@@
Transformer models are also integrated with GAN models for improved performance. Zhang et al. [173] use Relative GAN (Relativistic GAN) along with transformer modules. The Relative GANs enhance the discriminator by predicting the probability that a given sample is more realistic than other samples, rather than directly classifying a sample as real or fake. In this approach, the generator is not only responsible for producing realistic samples but also for maintaining the relative authenticity differences between samples. The relativistic discriminator improves training by reducing mode collapse and promoting diversity in generated outputs. Additionally, the internal structure of their model incorporates transformer modules to effectively capture temporal relationships, further enhancing the generation of time-dependent data, such as network traffic.
The authors in [132] explored the integration of autoencoders and LSTMs with GANs to enhance generation quality, particularly for improving NIDS. In their framework, an autoencoder was first trained to learn a latent representation of real sequences of packet sizes. This autoencoder consists of LSTM layers to capture temporal dependencies in sequential data, particularly the ordering of the packet sizes in a sequence. Subsequently, GAN was trained on this latent space to generate latent vectors, which could then be decoded into realistic sequences of packet sizes. In a similar approach,  [123] propose a GAN model combined with a LSTM approach for intrusion detection in fog computing network environments. After collecting network data, noise is removed, and features are filtered before generating data using the GAN model. This processed data is then used to train an LSTM model for detecting anomalous traffic in fog computing networks.
V-B5 Miscellaneous approaches
@@
The authors in [45] propose a model that combines multiple GAN architectures—Graph GAN, Image GAN, and Sequence GAN—to process different data types in parallel, generated by style-based encoders. Graph GAN processes graph data that encodes relationships between packet IP addresses. Image GAN takes image data created by converting time intervals into visual representations. Sequence GAN handles packet-layer information (e.g., Ethernet, IP) as numerical sequences. Additionally, they generate optional data, such as packet loads, using Sequence GAN and Markov models. The synthetic data produced by these trained models is decoded to recreate .pcap files.
Yang et al. [168] focus on generating network traffic using traffic obfuscation techniques. They use Gaussian Mixture Models (GMM) for data normalization and autoencoders to reduce network data to a low-dimensional form. This processed data is then fed into a GAN model that integrates Swing Transformers. Swing Transformers are a specialized type of transformer architecture designed to capture feature variations at multiple scales within deep learning layers. The use of Swing transformers has enabled the model to capture feature variations across different scales within the deep learning layers while also integrating self-attention mechanisms.
@@
Khan et al.[70] proposed a data generation approach for SDN settings in healthcare consumer IoT device management, specifically addressing class imbalance. The method utilizes a Boundary Equilibrium Generative Adversarial Network (BEGAN), with a separate GAN model dedicated to each data class. The training process is halted based on a convergence metric to ensure optimal performance. Additionally, autoencoders are employed for dimensionality reduction in the data, enhancing the efficiency of the model.
In addition to the practical data generation, a thorough theoretical analysis of GAN-based data generation is studied. For example, [88] explores the privacy implications of synthetic data generated by GANs, focusing on differential privacy and membership inference attack resilience. The authors prove that generated samples inherently have weak differential privacy properties, scaling with the ratio of generated to training samples. They also show that vanilla GANs, without special privacy mechanisms, are robust against black-box membership inference attacks. Their contributions lie in providing theoretical bounds on privacy risks for GAN-generated samples, highlighting areas where traditional GANs meet privacy requirements and where further techniques are needed.
V-B6 Limitations in GAN based traffic generation
@@
V-B7 Summary
GANs are among the most widely used generative AI models for network traffic generation. Table V provides a comprehensive summary of the work done with GANs. To optimize these models for traffic generation, various GAN variants have been proposed, particularly by modifying their internal neural network architectures. These adaptations include integrating autoencoders, recurrent neural networks (RNNs), and incorporating conditional information. Such enhancements have successfully facilitated traffic generation for a wide range of network-related applications, including intrusion detection and improving other machine learning tasks, primarily for both packet-level and flow-level data. While GANs inherently face challenges such as mode collapse and training instability, existing approaches have effectively addressed these issues through advanced mechanisms, enabling the generation of high-fidelity data.
V-C Diffusion Models (DM)
@@
DDPM Sivaroopan et al. [136] Anomaly detection, real-time classification, traffic fingerprinting DDPM Transport, Network Packet level Yes
Chi et al. [31] Generic RF signal generation Hierarchical Diffusion Transformer Physical Raw signal data Yes
Stable DM Jiang et al. [64, 65] Generic network traffic generation Stable DM + ControlNet [178] Transport, Network Packet level Yes
Controlled DM Zhang et al. [176] Mobile app usage modeling and service interaction tracing Hierarchical DM + Word2Vec + Encoder-Decoder + Two-layer transformer Transport, Application Network flow traces, App usage sequences No
Chai et al. [25] Network optimization DM + Attention-based classifier-free guidance + Contrastive learning Application Aggregated mobile traffic data across geographic areas Yes
Chai et al. [26] Mobile traffic generation, Network optimization Conditional DM + Knowledge graph + Frequency attention mechanism Application Mobile traffic time-series, Urban data (POI, population, regional functions) No
Qi et al.[122] 5G network traffic generation CANDLE + GCN Application Traffic time-series, Geospatial grid-based traffic patterns No
Jasni et al. [62] Adversarial packet generation, Evasion of NIDS DM + BERT Transport, Network Flow level No
Lightweight DM Li et al. [83] Synthesis of malicious traffic for edge devices Lightweight DM with depthwise separable convolutions Network Packet level No
DMs were first introduced to the deep learning community in 2015 by [140], where they were initially conceptualized as probabilistic generative models. However, it was not until 2020, with the groundbreaking work in [60], that DMs gained significant traction. The development and training of these models in this period marked a turning point in generative modeling, with DMs quickly emerging as a compelling alternative to long-dominant GANs.
@@
V-C1 Denoising Diffusion Probabilistic Model (DDPM)
@@
Generating high-fidelity RF signals presents unique challenges due to the need to capture their full structure across the time, frequency, and complex value domains. In  [31], the authors highlight that RF signals, as time-series data, capture dynamic details that standard diffusion models struggle to synthesize. Moreover, critical RF features embedded in the frequency domain, such as Doppler shifts, are often overlooked due to the focus of DMs on spatial and amplitude information. Furthermore, the complex-valued nature of RF signals, including phase data, is neglected in existing models. Addressing these complexities, the authors introduce RF-Diffusion to synthesize RF signals utilizing a novel Time-Frequency Diffusion (TFD) theory and a Hierarchical Diffusion Transformer (HDT). In TFD theory, they present RF-Diffusion model which alternates between adding noise in the time domain and blurring the frequency domain during the forward process, and reverses these effects by denoising and deblurring. This restoration emphasizes the accuracy of the amplitude in the time domain and the continuity of the frequency domain. The HDT incorporates a hierarchical architecture to decouple spatio-temporal dimensions, attention-based diffusion blocks with enhanced Transformers for feature extraction, and a complex-valued design to encode both amplitude and phase information. The authors synthesize Wi-Fi and Frequency-Modulated Continuous Wave (FMCW) signals using RF-Diffusion which outperforms typical DDPM, DCGAN, and CVAE-based methods.
V-C2 Stable DM
@@
V-C3 Controlled DM
@@
Expanding the scope of diffusion models in network traffic generation and cybersecurity, recent research has focused on improving controllability, reducing data dependency, and capturing complex spatial-temporal patterns - key areas where traditional GAN-based approaches often fall short. Zhang et al.  [176] presented a hierarchical, service-guided, controlled diffusion model for network flow trace generation. This approach was designed to explicitly model users’ dynamic network usage patterns, focusing on the interactions between users and network services, such as mobile application usage traces. Their hierarchical generation structure consisted of multiple layers, the first of which was responsible for modeling app usage behavior. This app usage data was then used to guide the generation of network flow traces, ensuring that the generated flows reflected realistic service usage patterns. The model captured co-usage relationships between different services, as well as sequential usage patterns, by leveraging a pre-trained embedding model and an encoder-decoder architecture. Furthermore, it incorporated a two-layer transformer network to model the temporal and feature correlations inherent in multidimensional network flow data. The generation process followed a two-step structure: first, a diffusion model generated the app usage patterns, and subsequently, a second diffusion model was trained on network flow traces, using the app usage patterns to guide the generation of realistic network traffic data.
@@
In [25], the authors introduce OpenDiff, a model that leverages publicly available data such as satellite imagery, POI distributions, and population density to generate mobile traffic data for network planning and optimization. By adopting a multi-positive contrastive learning algorithm and an attention-based classifier-free guidance module, OpenDiff mitigates the reliance on non-public data and enhances the controllability of traffic generation. Data pre-processing steps included feature extraction from satellite images and POI, population distribution analysis, and multi-positive sample construction for contrastive learning.
@@
The paper, CANDLE [122], uses a conditional diffusion model to simulate 5G network traffic in regions with limited historical data, conditioning on existing 4G traffic patterns. This model incorporates graph convolutional networks (GCN) and a cross-attention mechanism to address data scarcity and provide operators with insights for base station deployment. In addition, in [26], KG-Diff is proposed to incorporate urban contextual features from an Urban Knowledge Graph (UKG) to guide mobile traffic generation. By employing a frequency attention mechanism that decomposes network traffic into multiple periodic components, KG-Diff effectively models complex temporal patterns and the underlying correlations between urban environments and network traffic, offering a more nuanced approach to mobile traffic generation. Fourier transform was used for frequency decomposition.
@@
Extending the application of DMs to cybersecurity, DiffuPac [62] innovatively combines a Bidirectional Encoder Representations of Transformers (BERT) model with a diffusion framework to generate adversarial network packets capable of evading detection by NIDS. DiffuPac introduces a targeted noising mechanism and classifier-free guidance, ensuring that only malicious packet sequences are noised while preserving the integrity of normal traffic. Additionally, the model employs a unique packet concatenation strategy to seamlessly integrate malicious packets into benign traffic flows, leveraging contextual embeddings from BERT for enhanced mimicry. DiffuPac’s fine-tuning phase optimizes a variational lower bound that prioritizes the reconstruction of malicious packets as indistinguishable from normal traffic patterns. Evaluations demonstrate that DiffuPac not only outperforms traditional approaches like GANs and LSTMs but also operates under the realistic constraint of limited attacker knowledge, showcasing its practical efficacy in real-world scenarios.
V-C4 Light-weight DMs

### v2

@@ IV-B1 Vanilla GAN
Advancing the usage of vanilla GAN models, Xiao et al.[184] proposed a federated learning(FL)-based framework for distributed GAN-based network traffic generation with self-supervised learning. The approach involves local model training at the edge, where each edge node contains multiple generators, a discriminator, and a classifier. Each generator is trained on a local dataset, while the discriminator helps improve the performance of all generators at the edge. Locally trained discriminators and generators from all edge nodes are then aggregated globally using the Federated Averaging (FedAvg) algorithm. The classifier at each edge provides pseudo-labels, enabling self-supervised training of the GAN models and facilitating autonomous training of the entire system.
@@
In other network applications, vanilla GANs are utilized to generate network traffic data to enhance network IDS, particularly when class imbalance is prevalent in anomalous data. In this effort, Shahriar et al. [153] apply an encoding approach to convert categorical input features into a low-dimensional continuous space. They also employ feature normalization techniques and reduce feature dimensionality using Principal Component Analysis (PCA) to support the GAN model training process.
IV-B2 Wasserstein GAN (WGAN)
@@
In [55], the authors leverage WGAN to obfuscate network traffic, reducing the detection capabilities of various applications. Initially, they derive statistical features for the flow level data they collected followed by normalization across the flows. Then, they use WGAN to collect network traffic data and generate flow-level statistics, followed by applying WGAN again to synthesize network packet traces based on these synthetic flow-level statistics to produce obfuscated traffic. The model has multiple WGAN modules trained on different network traffic based on applications (e.g., Google Chrome, Google Drive), enabling application-level traffic obfuscation.
@@
Researchers have applied conversion of network traffic data into the image domain, where WGANs show great promise, after making multiple modifications to improve GANs for 1D data generation. In [157], the authors employ the Gramian Angular Summation Field (GASF) imaging mechanism to transform certain attributes of network traces into visual representations [179]. These images are utilized for network traffic generation through a simple WGAN framework. This imaging approach effectively captures the correlations between time series points, which previous studies attempt to model using more complex GAN architectures incorporating LSTMs and other networks. The authors develop a post-processing mechanism for the generated images and evaluate the model across multiple traffic types, demonstrating improved performance with the straightforward WGAN for image generation.
WGANs are used in RF signal synthesis. However, training GANs on long time-domain sequences of RF signals is challenging. To address this, [150] proposed representing RF signals in the spectral domain, allowing the use of existing GAN architectures. The authors consider 2.5s long 4G LTE signal samples at 40 MSamples/sec which were transformed into 256x256 compact spectrograms, where the horizontal axis represented the time dimension over approximately 2 ms, and the vertical axis represented the frequency offset within a 40 MHz band. Each spectrogram served as input to the GAN, effectively framing the RF signal generation as a 2D problem similar to image synthesis. The authors also experimented with direct training of GANs on time series data. They found that while this approach could generate raw I/Q samples, it was significantly more challenging to capture the fine-grained features and multiscale details of LTE signals compared to the spectral domain representation. Further, in [16], the authors propose a WGAN-based approach for estimating frequency-selective channels at high frequencies (e.g., mmWave, THz) using few pilots and operating under low SNR conditions. The generative model is first trained to capture the true channel distribution and is then used as a prior to estimate current channels by optimizing its input based on received signals.
IV-B3 Conditional GAN (CGAN)
@@
In [54], the authors propose Auxiliary CGAN (ACGAN) that extends adversarial learning to generate synthetic radar time-frequency domain signatures for human activity recognition. This approach uses Time-Frequency images for data synthesis. In [50], the authors propose an unsupervised over-the-air (OTA) algorithm that trains CGAN with federated training setup using only noisy pilot measurements to estimate high-dimensional beamspace MIMO channels. GANSec [187], a CGAN framework, is proposed for augmenting wireless time-series data to improve anomaly detection under data scarcity and class imbalance. They explore various architectures and training objectives, demonstrating that detectors trained solely on GANSec-generated data perform well in cross-scenario jamming detection using real-world 5G measurements.
IV-B4 GANs with sequential models
@@
Transformer models are also integrated with GAN models for improved performance. Zhang et al. [200] use Relative GAN (Relativistic GAN) along with transformer modules. The Relative GANs enhance the discriminator by predicting the probability that a given sample is more realistic than other samples, rather than directly classifying a sample as real or fake. In this approach, the generator is not only responsible for producing realistic samples but also for maintaining the relative authenticity differences between samples. The relativistic discriminator improves training by reducing mode collapse and promoting diversity in generated outputs. Additionally, the internal structure of their model incorporates transformer modules to effectively capture temporal relationships, enhancing the generation of time-dependent data (e.g., network traffic).
The authors in [152] explored the integration of autoencoders and LSTMs with GANs to enhance generation quality, particularly for improving Network IDS (NIDS). In their framework, an autoencoder was first trained to learn a latent representation of real sequences of packet sizes. This autoencoder consists of LSTM layers to capture temporal dependencies in sequential data, particularly the ordering of the packet sizes in a sequence. Subsequently, GAN was trained on this latent space to generate latent vectors, which could then be decoded into realistic sequences of packet sizes. In a similar approach, [142] propose a GAN model combined with a LSTM approach for IDS in fog computing network environments. After collecting network data, noise is removed, and features are filtered before generating data using the GAN model. This processed data is then used to train LSTM for detecting anomalous traffic in fog computing networks.
IV-B5 Miscellaneous approaches
@@
The authors in [51] propose a model that combines multiple GAN architectures—Graph GAN, Image GAN, and Sequence GAN—to process different data types in parallel, generated by style-based encoders. Graph GAN processes graph data that encodes relationships between packet IP addresses. Image GAN takes image data created by converting time intervals into visual representations. Sequence GAN handles packet-layer information (e.g., Ethernet, IP) as numerical sequences. Additionally, they generate optional data, such as packet loads, using Sequence GAN and MMs. The synthetic data produced by these trained models is decoded to recreate .pcap files.
Yang et al. [191] generate network traffic using traffic obfuscation techniques. They use GMM for data normalization and autoencoders to reduce network data to a low-dimensional form. This processed data is then fed into a GAN model that integrates Swing Transformers. Swing Transformers are a specialized type of transformer architecture designed to capture feature variations at multiple scales within DL layers. The use of Swing transformers has enabled the model to capture feature variations across different scales within the DL layers while also integrating self-attention mechanisms.
@@
Khan et al.[84] proposed a data generation approach for Software-Defined Networking (SDN) settings in healthcare consumer IoT device management, specifically addressing class imbalance. The method utilizes a Boundary Equilibrium GAN (BEGAN), with a separate GAN model dedicated to each data class. The training process is halted based on a convergence metric to ensure optimal performance. Additionally, autoencoders are employed for dimensionality reduction in the data, enhancing the efficiency of the model.
In addition to the practical data generation, theoretical analysis of GAN-based data generation is also studied. For example, [103] explores the privacy implications of synthetic data generated by GANs, focusing on differential privacy and membership inference attack resilience. The authors prove that generated samples inherently have weak differential privacy properties, scaling with the ratio of generated to training samples. They also show that vanilla GANs, without special privacy mechanisms, are robust against black-box membership inference attacks. Their contributions lie in providing theoretical bounds on privacy risks for GAN-generated samples, highlighting areas where traditional GANs meet privacy requirements and where further techniques are needed.
IV-B6 Limitations in GAN based traffic generation
@@
IV-B7 Summary
GANs are among the most widely used generative AI models for network traffic generation. Table V provides a comprehensive summary of the work done with GANs. To optimize these models for traffic generation, various GAN variants have been proposed, particularly by modifying their internal neural network architectures. These adaptations include integrating autoencoders, RNNs, and incorporating conditional information. Such enhancements have successfully facilitated traffic generation for a wide range of network-related applications, including IDS and improving other machine learning tasks, primarily for both packet-level and flow-level data. While GANs inherently face challenges such as mode collapse and training instability, existing approaches have effectively addressed these issues through advanced mechanisms, enabling the generation of high-fidelity data.
IV-C Diffusion Models (DM)
@@
DDPM [156] Anomaly detection, traffic fingerprinting DDPM Transport, Network Packet level Yes
[35] Generic RF signal generation HDT Physical Raw signal data Yes
[175] Privacy against MIA RF-Diffusion Physical WI-FI signal No
[126] RF data generation DDIM Physical RF signal No
[196] RFFI for IoT device authentication DDPM Physical RF signal No
[199] Radar signal synthesis in intelligent transportation system DM + SMOTE Physical Radar signal No
Stable DM [76, 77] Generic network traffic generation Stable DM + ControlNet [205] Transport, Network Packet level Yes
[178] High-resolution mmWave imaging Stable DM Physical mmWave No
Controlled DM [203] Mobile app usage modeling and service interaction tracing Hierarchical DM + Word2Vec + Encoder-Decoder + Two-layer transformer Transport, Application Network flow traces, App usage sequences No
[29] Network optimization DM + Attention-based classifier-free guidance + Contrastive learning Application Aggregated mobile traffic data across geographic areas Yes
[30] Mobile traffic generation, Network optimization Conditional DM + KG + Frequency attention Application Mobile traffic time-series, Urban data No
[141] 5G network traffic generation CANDLE + GCN Application Traffic time-series, Geospatial grid-based traffic patterns No
[74] Adversarial packet generation, Evasion of NIDS DM + BERT Transport, Network Flow level No
[172] Secured sensing system Conditional DM Physical CSI data No
[58] Denoising AE signals in concrete Conditional DM + CNN + Attention Physical AE signal No
[34] Depth estimation Conditional DM Physical CSI data Yes
[173] ISAC Conditional DM Physical CSI data Yes
Lightweight DM [97] Synthesis of malicious traffic for edge devices DM with depthwise separable convolutions Network Packet level No
DMs were first introduced to the DL community in 2015 by [159], where they were initially conceptualized as probabilistic generative models. However, it was not until 2020, with the groundbreaking work in [70], that DMs gained significant traction. The development of these models marked a turning point in generative modeling, with DMs quickly emerging as a compelling alternative to long-dominant GANs.
@@
IV-C1 Denoising Diffusion Probabilistic Model (DDPM)
@@
Generating high-fidelity RF signals presents unique challenges due to the need to capture their full structure across the time, frequency, and complex value domains. In [35], the authors highlight that RF signals, as time-series data, capture dynamic details that standard diffusion models struggle to synthesize. Moreover, critical RF features embedded in the frequency domain, such as Doppler shifts, are often overlooked due to the focus of DMs on spatial and amplitude information. Furthermore, the complex-valued nature of RF signals, including phase data, is neglected in existing models. Addressing these complexities, the authors introduce RF-Diffusion to synthesize RF signals utilizing a novel Time-Frequency Diffusion (TFD) theory and a Hierarchical Diffusion Transformer (HDT). In TFD theory, they present RF-Diffusion model which alternates between adding noise in the time domain and blurring the frequency domain during the forward process and reverses these effects by denoising and deblurring. This restoration emphasizes the accuracy of the amplitude in the time domain and the continuity of the frequency domain. The HDT incorporates a hierarchical architecture to decouple spatio-temporal dimensions, attention-based diffusion blocks with enhanced transformers for feature extraction, and a complex-valued design to encode both amplitude and phase information. The authors synthesize Wi-Fi and FMCW signals using RF-Diffusion which outperforms typical DDPM, DCGAN, and CVAE-based methods.
Building on top of the RF-diffusion model, authors of [175] added differential privacy (DP) to the base model for privacy-preserving WI-FI data generation.They had proposed a hybrid training method for the DM applied to wireless data as a defense against Membership Inference Attacks (MIA). The approach involves initially training the model without privacy constraints. After a specified number of training rounds, DP is incorporated for fine-tuning. During this second phase, a cooptimization process is conducted in parallel to counteract the effects of the added noise.
In [126], the authors propose a DM tailored for RF data generation using Denoising Diffusion Implicit Models (DDIM) with noise schedules optimized for In-phase/Quadrature (I/Q) signal types. Unlike DDPM, this approach reduces computational overhead while maintaining high-quality output. In [196], the authors leverage DM to enhance RF Fingerprint Identification (RFFI) for IoT device authentication under low SNR conditions. By training a robust noise predictor and designing a noise removal algorithm, the approach effectively denoises received signals, restoring subtle hardware features crucial for device identification. This was demonstrated on Wi-Fi signal. In [199], the authors propose SMOTE-Diffusion, a DM combined with the SMOTE to improve stationary object classification using Impulse Radio UltraWideBand (IR-UWB) radar. While FMCW radar excels at detecting moving objects, IR-UWB offers high temporal resolution suitable for stationary object recognition but suffers from limited time-domain data. SMOTE-Diffusion addresses this by generating diverse, realistic synthetic radar signals, reducing data collection needs and enhancing model performance.
IV-C2 Stable DM
@@
In [178], the authors propose a diffusion-based neural network for high-resolution mmWave imaging, which transforms mmWave signals into high-quality images by leveraging the intrinsic features of target scenes. This approach uses conditional generation through stable diffusion to enhance image quality without relying on sparsity assumptions. To support this, they design an optimized mmWave metasurface and codebook that improve signal penetration and phase control, enabling effective imaging with compact, off-the-shelf hardware and no mechanical movement.
IV-C3 Controlled DM
@@
Expanding the scope of diffusion models in network traffic generation and cybersecurity, recent research has focused on improving controllability, reducing data dependency, and capturing complex spatial-temporal patterns – key areas where traditional GAN-based approaches often fall short. Zhang et al. [203] presented a hierarchical, service-guided, controlled diffusion model for network flow trace generation. This approach was designed to explicitly model users’ dynamic network usage patterns, focusing on the interactions between users and network services, such as mobile application usage traces. Their hierarchical generation structure consisted of multiple layers, the first of which was responsible for modelling app usage behaviour. This app usage data was then used to guide the generation of network flow traces, ensuring that the generated flows reflected realistic service usage patterns. The model captured co-usage relationships between different services, as well as sequential usage patterns, by leveraging a pre-trained embedding model and an encoder-decoder architecture. Furthermore, it incorporated a two-layer transformer network to model the temporal and feature correlations inherent in multidimensional network flow data. The generation process followed a two-step structure: first, a diffusion model generated the app usage patterns, and subsequently, a second diffusion model was trained on network flow traces, using the app usage patterns to guide the generation of realistic network traffic data.
@@
In [29], the authors introduce OpenDiff (i.e., for satellite image data), Point of Interest (POI) distributions, and population density to generate mobile traffic data for network planning and optimisation. By adopting a multi-positive contrastive learning algorithm and an attention-based classifier-free guidance module, OpenDiff mitigates the reliance on non-public data and enhances the controllability of traffic generation. Data pre-processing steps included feature extraction from satellite images and POI, population distribution analysis, and multi-positive sample construction for contrastive learning.
@@
The paper, CANDLE [141], uses a conditional diffusion model to simulate 5G network traffic in regions with limited historical data, conditioning on existing 4G traffic patterns. This model incorporates graph convolutional networks (GCN) and a cross-attention mechanism to address data scarcity and provide operators with insights for base station deployment. In addition, in [30], KG-Diff is proposed to incorporate urban contextual features from an Urban Knowledge Graph (UKG) to guide mobile traffic generation. By employing a frequency attention mechanism that decomposes network traffic into multiple periodic components, KG-Diff effectively models complex temporal patterns and the underlying correlations between urban environments and network traffic, offering a more nuanced approach to mobile traffic generation. The Fourier transform was used for frequency decomposition.
@@
Extending the application of DMs to cybersecurity, DiffuPac [74] innovatively combines a Bidirectional Encoder Representations of Transformers (BERT) model with a diffusion framework to generate adversarial network packets capable of evading detection by NIDS. DiffuPac introduces a targeted noising mechanism and classifier-free guidance, ensuring that only malicious packet sequences are noised while preserving the integrity of normal traffic. Additionally, the model employs a unique packet concatenation strategy to seamlessly integrate malicious packets into benign traffic flows, leveraging contextual embeddings from BERT for enhanced mimicry. DiffuPac’s fine-tuning phase optimises a variational lower bound that prioritises the reconstruction of malicious packets as indistinguishable from normal traffic patterns. Evaluations demonstrate that DiffuPac not only outperforms traditional approaches like GANs and LSTMs but also operates under the realistic constraint of limited attacker knowledge, showcasing its practical efficacy in real-world scenarios.
In [172], a DM-based secure sensing system is proposed to enhance the security of integrated sensing and communications by preventing unauthorized sensing using CSI. The system uses discrete and continuous conditional diffusion models to selectively activate network elements and generate safeguarding signals that mask user activity. This ensures that only authorized devices can recover accurate CSI for sensing. AEDM [58], a DM-based DL approach for denoising acoustic emission (AE) signals in concrete, aimed at improving structural health monitoring. The model combines a CNN with attention mechanisms and multi-scale convolutions, and conditions the diffusion process on noise observations to effectively separate noise from relevant signals. This enables accurate damage detection despite challenges like material anisotropy and external interference. WiViD [34], a diffusion-based depth estimation system that combines commercial Wi-Fi signals with vision to overcome the limitations of Light Detection and Ranging (LiDAR), mmWave, and monocular vision methods. By introducing a Multimodal Conditional Diffusion (MMCD) mechanism, WiViD iteratively refines depth predictions for improved accuracy. The system features two key encoders—the Complex-Valued CSI Encoder (CCE) and the Residual Image Encoder (RIE)—to effectively capture and fuse spatio-temporal features from Wi-Fi CSI and RGB images. This fusion enables robust and high-precision depth estimation suitable for real-world applications. In [173], the authors propose a DM-based data augmentation system to address the challenge of limited CSI data in Integrated Sensing and Communication (ISAC) systems. The approach trains a conditional DM on real-world samples to generate additional CSI data, increasing sample quantity. A second DM is then applied to denoise the generated data to improve quality. This method supports more effective AI-driven CSI analysis for physical space and human activity monitoring in 6G networks.
IV-C4 Light-weight DMs